# Through the Magnifying Glass: Adaptive Perception Magnification for Hallucination-Free VLM Decoding

### v2

@@ Abstract
Existing vision-language models (VLMs) often suffer from visual hallucination, where the generated responses contain inaccuracies that are not grounded in the visual input. Efforts to address this issue without model finetuning primarily mitigate hallucination by reducing language biases reducing bias contrastively or amplifying the weights of visual embedding during decoding. However, these approaches improve visual perception at the cost of impairing the language reasoning capability. In this work, we propose the Perception Magnifier (PM), a novel visual decoding method that iteratively isolates relevant visual tokens based on attention and magnifies the corresponding regions, spurring the model to concentrate on fine-grained visual details during decoding. Specifically, by magnifying critical regions while preserving the structural and contextual information at each decoding step, PM allows the VLM to enhance its scrutiny of the visual input, hence producing more accurate and faithful responses. Extensive experimental results demonstrate that PM not only achieves superior hallucination mitigation but also enhances language generation while preserving strong reasoning capabilities. Code is available at https://github.com/ShunqiM/PM.
1 Introduction
The evolution of large vision-language models (VLMs) [40, 67, 61, 2, 70] has revolutionized the integration of computer vision and natural language processing, enabling seamless interaction between visual and textual modalities. These models effectively translate complex visual inputs into coherent and contextually relevant textual descriptions, driving advancements across a wide range of applications [65, 17, 60]. Fueled by innovations in model architectures [14, 18], training methodologies [50, 40, 12], and diverse multimodal datasets [53, 54], large VLMs continue to expand their capabilities and adaptability, solidifying their role as foundational tools in modern AI systems.
Despite their outstanding abilities, VLMs are prone to hallucination, where generated outputs fail to align with the visual input, with inaccuracies or unfaithful content. While some hallucinations stem from VLMs adopting the same architecture as pretrained large language models (LLMs) and inheriting their intrinsic limitations [42], the integration of visual and language features can also lead to over-reliance on language priors or limited visual representation capacity.
@@ 1 Introduction
Various approaches have been proposed to mitigate hallucination, ranging from training-end techniques, such as debiased dataset [39, 77] or enlarged visual resolution [75, 38], to inference end methods such as post-processing for hallucination removal [72, 82] and specialized decoding techniques [36, 24, 34]. Among these, hallucination-mitigation decoding methods have gained significant attention for their ability to improve visual faithfulness without requiring additional finetuning of the model. These methods primarily aim to enhance the contribution of visual information during model reasoning to mitigate hallucination. This enhancement can be achieved explicitly by increasing the weight of visual embeddings [43, 84], or implicitly by contrastively penalizing outputs that exhibit vague [34, 33], absent [79, 19], or confused visual interpretations [62, 5]. However, these methods face limitations in scenarios where insufficient visual resolution causes object misinterpretation. While a concurrent study [76] explores resizing a cropped key visual region as an additional input, it erases structural and contextual information from the image and confuses the VLM by giving multiple input images.
Inspired by the observation that VLMs allocate only marginal attention to many visual regions during inference, we propose the Perception Magnifier (PM), a visual decoding approach that adaptively modulates spatial resolution, processing key visual regions at higher detail while compressing less relevant areas, as illustrated in Fig. 1. By refining the visual input during decoding, PM enhances attention to fine-grained details without compromising structural visual context of the image or the intrinsic reasoning ability of VLM. Unlike methods that apply static modifications to input images [73, 68, 69, 76], PM dynamically adjusts the image representation throughout decoding, enabling more effective open-ended generation.
Specifically, we introduce a gradient-free approach to iteratively identify high-attention visual tokens relevant to generation across deep layers and aggregate their attention scores into a token-level heatmap. This heatmap is post-processed to enhance coverage and smoothness, then upsampled to construct a pixel-level perception map. Guided by this map, PM refines the visual input by magnifying relevant regions while compressing less critical areas, implemented as a resampling process [80, 27]. This dynamic adjustment allows the model to iteratively refine its focus throughout decoding, ensuring improved visual faithfulness without disrupting language reasoning.
@@
2 Related Work
2.1 Vision-Language Models
The scaling up of computational resources and the availability of large-scale data has driven advances in pre-training [49, 15, 3] and instruction tuning techniques [48, 64, 45], allowing the emergence of numerous large language models (LLMs) [58, 81, 3, 46, 10] that can generate human-like text and perform complex reasoning across diverse tasks. Based on the advancements in LLMs, many vision-language models (VLMs) have been developed by integrating powerful language models with vision encoders using connection modules. While both cross-attention mechanisms [35, 12, 1, 83] and projector-based token concatenation [40, 2, 61, 59, 32, 60] were explored in earlier works, modern models [41, 22, 67, 6, 17] usually adopt the latter for its simplicity and efficiency in aligning visual and textual features. Although these VLMs leverage the power of existing LLMs by aligning image features with the text modality, their autoregressive nature and the limitations of visual model capabilities still lead to hallucination issues.
2.2 Hallucination in Vision-Language Models
In addition to hallucination inherited from the LLM backbones [42] due to insufficient context attention and capability misalignment [23, 78, 26], VLMs also face challenges unique to their multimodal nature, such as language bias [84, 79], where the language decoder may rely heavily on preceding textual tokens over visual tokens. To address this, various methods have been proposed to enhance the role of visual information by prompting input images [73, 68, 69] or amplifying the influence of visual tokens [79, 84, 43]. VLMs also hallucinate due to the limitations of vision encoders or connection modules [42]. Prior works have addressed this by increasing visual tokenization resolution [75, 38] or improving the capacity of these modules [41, 8]. Efforts to control hallucination have also focused on the data aspect, including mitigating bias [31, 39, 77, 56, 47, 74], creating biased samples for contrastive learning [28], and developing methods to evaluate hallucination [30, 21, 71, 63], supported by various metrics and algorithms for quantifying hallucination [52, 37, 44, 29, 7]. Post-processing techniques have also been proposed to eliminate hallucinations by identifying visually ungrounded outputs [72, 82]. While these approaches tackle hallucination from multiple angles, few methods focus on enhancing visual recognition capabilities beyond increasing resolution. Instead, we propose a novel approach that assists VLMs with a magnifier to capture fine-grained details in input images effectively. A concurrent but slightly earlier study, ViCrop [76], adopts a similar idea by cropping and resizing a key visual bounding box based on attention maps as additional inputs. However, it does destruct the integrity of image structure, and can fail when spatial awareness is required, whereas our method preserves the visual input structure by magnifying key regions while compressing less relevant areas.
2.3 Hallucination Mitigation Decoding Strategies
Decoding strategies determine how a model generates the next text token, with methods like beam search [57] enabling higher-quality text generation without additional fine-tuning. In the context of hallucination control, contrastive decoding (CD) [36] is a popular approach that optimizes output logits by amplifying patterns favored by larger expert LMs while suppressing a biased output more prone to hallucination, typically derived from smaller amateur LMs. DoLa [11] introduced an efficient CD formulation by treating early-exit logits of expert LMs as the biased output. Beyond CD, OPERA [24] mitigates hallucination in LMs by penalizing beam outputs that over-rely on specific aggregation tokens. For VLMs, VCD [34] extends CD by defining biased outputs as those generated with noisy image inputs. Subsequent works explored diverse biased output designs, such as logits derived without image inputs [79, 19], excluding key visual tokens [25], or producing hallucinative outputs via finetunning [5] or prompt-engineering [62]. Other methods aim to enhance expert outputs by increasing the weight of visual tokens in the model [43, 84] or searching for focal views with an external detector [9]. [33] further refines VCD by fusing outputs from multiple visually biased samples with an entropy-weighted strategy. Although these decoding methods effectively reduce hallucinations by penalizing less visually grounded outputs, treating them solely as ‘bias’ risks undermining the model’s ability to form logical inferences and maintain coherence in language-driven contexts. Instead, our method mitigates hallucination by adaptively magnifying relevant image regions, enhancing visual faithfulness while preserving the reasoning and language capabilities of VLMs.
3 Methods
@@
Figure 2: Structural diagram of the proposed PM method. At each decoding step, we aggregate multi-layer attentions into a token-level heatmap $\mathcal{H}^{(i)}$. By iteratively masking the most attended tokens and re-forwarding, we collect heatmaps $\{\mathcal{H}^{(i)}\}$ and combine them into $\mathcal{H^{*}}$ once the process ends, and post-processing $\mathcal{H^{*}}$ to obtain $\mathcal{H}_{proc}$ (yellow block). The magnifier resamples the visual input according to $\mathcal{H}_{proc}$, producing $\hat{I}$ (purple block). The VLM can then generate the next token with enhanced focus on critical regions with finer details.
In traditional autoregressive VLMs, the model takes an image $I$ and a text sequence $T^{i}$ at each decoding step. Given this input, it predicts the next token $T^{i+1}$ with: $O^{i+1}=VLM(I,T^{i}),$ where $O^{i+1}$ represents the output logits over the vocabulary. The model then applies the softmax function to obtain a probability distribution: $P(T^{i+1}\mid I,T^{i})=\text{softmax}(O^{i+1}).$ A sampling strategy is then applied to determine the final token. For example, in greedy decoding, the token with the highest probability is selected: $T^{i+1}=\arg\max P(T^{i+1}\mid I,T^{i}).$ This predicted token is appended to the sequence, and the process repeats until the end-of-sequence (EOS) token is generated.
@@
In this paper, we propose the Perception Magnifier (PM), a novel approach that iteratively identifies and enhances key visual regions to refine visual input for more effective token generation. As illustrated in Fig. 2, PM iteratively identifies highly attended visual tokens, blocks them, and re-forwards the sequence until no tokens receive significant attention. The identified tokens are then aggregated based on their attention scores to construct a perception map $\mathcal{P}$. This map is used to magnify the original image $I$ via a sampling process, $\hat{I}^{i}=\mathcal{T}(I,\mathcal{P})$. The refined image $\hat{I}^{i}$ then serves as the refined visual input for generating the next token, $T^{i+1}=\text{VLM}(\hat{I}^{i},T^{i})$, at the current decoding step. The following sections detail the perception map construction, iterative refinement for improving the coverage of perception map, and the attention-based magnification process in the proposed PM method.
3.1 Perception Map Construction
@@
In this section, we first introduce a simple method to construct a naive perception map to guide the magnifier. This method is applicable to any mainstream VLMs that adopt an interleaved architecture, where patchified visual embeddings are integrated into the input sequence alongside textual tokens. These models process information by interleaving system prompts, visual representations, and past textual interactions, such as user queries and generated responses, into a unified sequence for a backbone LM.
@@
Once the input sequence is formed, the transformer processes it via self-attention. For a single attention head, the attention weight from token $x_{i}$ to $x_{j}$ is computed as $\mathrm{Attn}_{i,:}=\text{Softmax}\left(Q_{i}K_{j}^{\top}/\sqrt{d_{k}}\right)$, where $Q_{i}$ and $K_{j}$ are the query and key vectors of tokens $x_{i}$ and $x_{j}$, respectively, and $d_{k}$ is the key dimension. The corresponding value $V_{j}$ is then weighted by $\mathrm{Attn}_{i,j}$ to obtain the hidden representation of $x_{i}$: $z_{i}=\sum_{j}\mathrm{Attn}_{i,j}\,V_{j}$. In practice, transformers stack multiple layers and heads, producing attention weights across different levels. These attention patterns can be represented as a multi-dimensional tensor: $\mathrm{Attn}\in\mathbb{R}^{N_{l},N_{h},H,W}$, where $N_{l}$ and $N_{h}$ denote the number of layers and attention heads, respectively, and $H\times W$ represents the spatial dimensions of visual tokens, which can be reshaped into a 2D feature map. PM utilizes the cached $\mathrm{Attn}$ to construct a token-level heatmap, highlighting the most relevant visual regions based on the model attention.
@@
As indicated by prior studies [13], deep networks primarily aggregate and interpret visual information in their earlier layers before leveraging these representations in later layers. Moreover, it has been observed that intermediate features can already yield competitive performance [11]. In line with these findings, we empirically note that middle-layer attention often more accurately localizes relevant objects than that of the final layers. Hence, we construct the perception map by aggregating self-attention matrices from layers $l\geq\mathcal{L}$ and applying a max pooling across all heads $h$ to highlight the impact of more visually important regions. Formally, the token-level heatmap $\mathcal{H}$ can be obtained by:
@@
The naive perception map constructed without iterative refinement could already be upsampled to guide the magnifier. Before up-sampling the token-level heatmap $\mathcal{H}$ to pixel-level perception map, we first post-process the heatmap with the method similar to [73] to obtain a more smooth and well-distributed attention distribution for better magnification quality. Firstly, we normalize $\mathcal{H}$ by scaling them to unit interval, obtaining $\mathcal{H}^{\prime}$. Then, we amplify the variance of the normalized map by a scaling coefficient of $\alpha$, and then compress its values back into the range $[0,1]$ using a sigmoid function, resulting in $\mathcal{H}_{\text{proc}}$:
@@
3.2 Iterative Refinement
While the naive perception map serves as a simple and effective guide for magnification, a single forward pass may fail to capture all instances of objects when multiple occurrences exist. This limitation arises because deep vision models exchange fine-grained features at lower layers and aggregate them into specific tokens that act as information registers [13, 51]. This aggregation concentrates visual information into a few key tokens, reducing the contribution of other visual tokens associated with semantically similar regions. As a result, when similar visual patterns appear in different parts of the image, VLMs may primarily attend to a subset of key tokens, leading to an incomplete perception map that overlooks certain regions of interest.
To enhance the coverage of perception map, we propose an iterative refinement approach on attention that compels VLMs to attend to additional visual regions by iteratively blocking key registers from their reasoning process. However, directly masking key visual regions at the image level would require fine-grained segmentation or ROI localization, making it computationally challenging. Moreover, such an approach is infeasible for decoding, as it would necessitate re-tokenization of visual tokens. Instead, we apply information blocking at the token level, effectively mitigating these challenges while preserving inference efficiency.
Specifically, at the first iteration, we forward the tokenized input through the VLM using the default attention mask $M^{(0)}$, initialized as a vector of ones, and obtain the initial token-level attention heatmap $\mathcal{H}^{(0)}$. To identify the most visually attended tokens $T^{(0)}$, we employ a clustering-based method that partitions the attention values into two clusters, selecting tokens in the one with the higher centroid as $T^{(0)}$, as visually important tokens typically exhibit significantly higher attention than others. To suppress their influence in subsequent iterations, we update the attention mask as: $M^{(1)}=M^{(0)}\cap\neg T^{(0)}$, where $\neg T^{(0)}$ represents the complement of the mask over $T^{(0)}$, ensuring that these highly attended tokens are excluded in the next pass. The updated input is then reprocessed by the VLM, producing a new token-level heatmap $\mathcal{H}^{(1)}$. This process is repeated iteratively until the total attention score in the next iteration falls below $\beta$ or an upper limits on iterations is reached, i.e., $\sum\mathcal{H}^{(i+1)}<\beta,\quad\text{for }0\leq i\leq 8,$ indicating that no additional visually relevant tokens remain to be allocated substantial amount of attention.
Finally, we construct the refined token-level heatmap $\mathcal{H}^{*}$ by aggregating attention across all $T$ iterations while normalizing for repeated contributions at each token position:
$\mathcal{H}^{*}=\frac{\sum_{i=0}^{T}\mathcal{H}^{(i)}}{\sum_{i=0}^{T}\mathbbm{1}_{M^{(i)}}}.$ (3)
The iteratively refined heatmap $\mathcal{H}^{*}$ is then post-processed to $\mathcal{H}_{proc}$ using the aforementioned method to produce a pixel-level perception map $\mathcal{P}$. This refined perception map captures all major visual elements relevant to next-token generation, providing an enhanced guide for the perception magnifier and ensuring focus on the most critical regions for the VLM.
3.3 Attention-Based Magnification
@@
Inspired by [80], we develop the perception magnifier as an attention-based sampling process. The perception map $\mathcal{P}$ is treated as a mass function, where regions with higher perception values are more likely to be sampled. To achieve this, we first decompose $\mathcal{P}$ into marginal distributions along the horizontal and vertical dimensions and compute their cumulative distributions $\mathcal{F}_{x}(n)$ and $\mathcal{F}_{y}(n)$ as follows:
$\begin{split}\mathcal{F}_{x}(n):=\sum_{j=1}^{n}\max_{1\leq i\leq w}\mathcal{P}_{i,j},\\ \mathcal{F}_{y}(n):=\sum_{i=1}^{n}\max_{1\leq j\leq h}\mathcal{P}_{i,j},\end{split}$ (4)

### v3

@@ Abstract
Existing vision-language models (VLMs) often suffer from visual hallucination, where the generated responses contain inaccuracies that are not grounded in the visual input. Efforts to address this issue without model finetuning primarily mitigate hallucination by contrastively reducing language biases or amplifying the weights of visual embedding during decoding. However, these approaches remain limited in their ability to capture fine-grained visual details. In this work, we propose the Perception Magnifier (PM), a novel visual decoding method that iteratively isolates relevant visual tokens based on attention and magnifies the corresponding regions, spurring the model to concentrate on fine-grained visual details during decoding. By magnifying critical regions while preserving the structural and contextual information at each decoding step, PM allows the VLM to enhance its scrutiny of the visual input, hence producing more accurate and faithful responses. Extensive experimental results demonstrate that PM not only achieves superior hallucination mitigation but also enhances language generation while preserving strong reasoning capabilities.
Introduction
The evolution of large vision-language models (VLMs) (Liu et al. 2023; Xue et al. 2024; Wang et al. 2023; Bai et al. 2023; Yang et al. 2023b) has revolutionized the integration of computer vision and natural language processing, enabling seamless interaction between visual and textual modalities. These models effectively translate complex visual inputs into coherent and contextually relevant textual descriptions, driving advancements across a wide range of applications (Wu et al. 2023; Driess et al. 2023; Wang et al. 2022). Fueled by innovations in model architectures (Desai and Johnson 2021; Eichenberg et al. 2022), training methodologies (Radford et al. 2021; Liu et al. 2023; Dai et al. 2023), and diverse multimodal datasets (Schuhmann et al. 2024; Sharma et al. 2018), large VLMs continue to expand their capabilities and adaptability, solidifying their role as foundational tools in modern AI systems.
Despite their outstanding abilities, VLMs are prone to hallucination, where generated outputs fail to align with the visual input, with inaccuracies or unfaithful content. While some hallucinations stem from VLMs adopting the same architecture as pretrained large language models (LLMs) and inheriting their intrinsic limitations (Liu et al. 2024c), the integration of visual and language features can also lead to hallucinations caused by over-reliance on language priors or limited visual representation capacity.
@@ Introduction
Various approaches have been proposed to mitigate hallucination, ranging from training-end techniques, such as debiased dataset (Liu et al. 2024a; Zhang et al. 2025b) or enlarged visual resolution (Zhai et al. 2023; Li et al. 2024), to inference end methods such as post-processing for hallucination removal (Yin et al. 2024; Zhou et al. 2024) and specialized decoding techniques (Li et al. 2023b; Huang et al. 2024; Leng et al. 2024). Among these, hallucination-mitigation decoding methods have gained significant attention for their ability to improve visual faithfulness without requiring additional finetuning of the model. These methods primarily aim to enhance the contribution of visual information during model inference to mitigate hallucination. This enhancement can be achieved explicitly by increasing the weight of visual embeddings (Liu, Zheng, and Chen 2025; Zhu et al. 2024), prompting key regions (Yu, Yu, and Wang 2024; Yang et al. 2024), or implicitly by contrastively penalizing outputs that exhibit vague (Leng et al. 2024; Lee, Tsai, and Chiu 2024), absent (Zhang et al. 2024; Favero et al. 2024), or confused visual interpretations (Wang et al. 2024a; Chen et al. 2024a). However, these methods face limitations in scenarios where insufficient visual resolution causes object misinterpretation.
Inspired by the observation that VLMs allocate marginal attention to many visual tokens during inference and that patch-based architectures struggle with scale-sensitive visual details, we propose the Perception Magnifier (PM), a visual decoding approach that adaptively modulates spatial resolution. As shown in Fig. 1, PM magnifies critical regions for higher-resolution processing while spatially compressing rather than discarding less relevant areas, thus enhancing fine-grained detail recognition without compromising structural visual context or intrinsic reasoning ability.
Specifically, we introduce a gradient-free approach to iteratively identify high-attention visual tokens relevant to generation across deep layers and aggregate their attention scores into a token-level heatmap. This heatmap is post-processed to enhance coverage and smoothness, then upsampled to construct a pixel-level perception map. Guided by this map, PM refines the visual input by magnifying relevant regions while compressing less critical areas, implemented as a resampling process (Zheng et al. 2019; Jia et al. 2022). This dynamic adjustment allows the model to iteratively refine its focus throughout decoding, ensuring improved visual faithfulness without disrupting language reasoning.
@@
Related Work
Hallucination in Vision-Language Models
Modern vision-language models (VLMs) typically connect vision encoders with large language models (LLMs) via either cross-attention mechanisms (Li et al. 2023a; Dai et al. 2023; Alayrac et al. 2022) or projector-based token concatenation (Liu et al. 2023; Bai et al. 2023; Koh, Salakhutdinov, and Fried 2023), with the latter dominating recent designs (Liu et al. 2024b; Hong et al. 2024; Xue et al. 2024) due to its simplicity and efficiency. While such designs align visual and textual modalities effectively, hallucination remains a persistent issue, arising not only from inherited LLM limitations (Liu et al. 2024c) but also from modality imbalance and vision encoder constraints.
To mitigate hallucination, prior efforts enhance visual grounding by prompting input images (Yu, Yu, and Wang 2024; Yang et al. 2024; Gizdov, Ullman, and Harari 2024), boosting visual token influence (Zhang et al. 2024; Zhu et al. 2024; Liu, Zheng, and Chen 2025), or improving vision module capacity (Zhai et al. 2023; Liu et al. 2024b; Chen et al. 2024b). Data-centric strategies include debiasing (Kim et al. 2025; Liu et al. 2024a; Zhang et al. 2025b; Ouali et al. 2025; Yue, Zhang, and Jin 2024), contrastive training (Jiang et al. 2024), and evaluation techniques (Kaul et al. 2024; Guan et al. 2024; Wang et al. 2024b; Li et al. 2023c; Lovenia et al. 2024; Jing et al. 2024), along with post-hoc filtering (Yin et al. 2024; Zhou et al. 2024). While these approaches tackle hallucination from multiple angles, few methods focus on enhancing visual recognition capabilities beyond increasing resolution. We propose a decoding strategy that magnifies semantically important regions without disrupting spatial coherence, improving perception during generation.
Hallucination Mitigation Inference Strategies
Some decoding-end approaches reduce hallucination by modifying next-token generation. Contrastive decoding (CD) (Li et al. 2023b) amplifies expert model logits while suppressing outputs from a biased source, typically a smaller or noisier model. Chuang et al. streamlines this by using early-exit logits from the expert model itself (2024). OPERA (Huang et al. 2024) penalizes beam outputs that heavily attend to generic aggregation tokens. VCD (Leng et al. 2024) adapts CD for VLMs by defining biased outputs from noisy image inputs. Later works extend this by deriving biased outputs by removing image inputs (Zhang et al. 2024; Favero et al. 2024) or enhance focus on textual feature (Wan et al. 2025), masking key visual tokens (Huo et al. 2024), or synthesizing biased outputs through finetuning (Chen et al. 2024a) or prompting (Wang et al. 2024a). Others enhance expert logits by upweighting visual tokens (Liu, Zheng, and Chen 2025; Zhu et al. 2024) or selecting focal views via external detectors (Chen et al. 2024c). Lee, Tsai, and Chiu fuse multiple biased outputs using weighted voting (2024). While these decoding strategies mitigate hallucination by penalizing visually inconsistent outputs, over-suppressing biased signals may hinder coherence and reasoning. Alternatively, some inference-time methods enhance visual grounding by applying masks or bounding boxes (Yu, Yu, and Wang 2024; Yang et al. 2024, 2023a), but still struggle to recognize fine-grained details. ViCrop (Zhang et al. 2025a) improves upon this by cropping and resizing key visual regions as auxiliary inputs. However, it disrupts spatial structure and may impair tasks requiring holistic understanding, and doubles visual tokens with more compute. In contrast, our method magnifies semantically important regions while compressing less relevant areas, preserving spatial integrity and supporting both visual grounding and coherent generation.
@@
Figure 2: Structural diagram of the proposed PM method. At each decoding step, we aggregate multi-layer attentions into a token-level heatmap $\mathcal{H}^{(i)}$. By iteratively masking the most attended tokens and re-forwarding, we collect heatmaps $\{\mathcal{H}^{(i)}\}$ and combine them into $\mathcal{H^{*}}$ once the process ends, and post-processing $\mathcal{H^{*}}$ to obtain $\mathcal{P}$ (yellow block). The magnifier resamples the visual input according to $\mathcal{P}$, producing $\hat{I}$ (purple block). The VLM can then generate the next token with enhanced focus on critical regions with finer details.
Methods
@@
In this paper, we propose the Perception Magnifier (PM), a novel approach that iteratively identifies and enhances key regions of the input image to refine visual input for less hallucinative token generation in VLMs. As illustrated in Fig. 2, PM iteratively identifies highly attended visual tokens, blocks them, and re-forwards the sequence until no tokens receive significant attention. The identified tokens are then aggregated based on their attention scores to construct a perception map $\mathcal{P}$. This map is used to magnify the original image $I$ via a sampling process, $\hat{I}^{i}=\mathcal{T}(I,\mathcal{P})$. The refined image $\hat{I}^{i}$ then serves as the refined visual input for generating the next-step output logit, $O^{i+1}=\text{VLM}(\hat{I}^{i},O^{i})$, where $O^{i}$ denotes the probability logits produced at step $i$. The following sections detail the perception map construction, the iterative refinement process for improving the coverage of perception map, and the attention-based magnification processed designed in the proposed PM method.
Perception Map Construction
@@
In this section, we first introduce a simple method to construct a basic perception map to guide the magnifier. This method applies to mainstream interleaved VLMs, where patchified visual embeddings are integrated with textual tokens into a unified input sequence processed by the VLM.
@@
The transformer processes the input sequence via self-attention. For a single attention head, the attention weight from token $x_{i}$ to $x_{j}$ is computed as $\mathrm{Attn}_{i,:}=\text{Softmax}\left(Q_{i}K_{j}^{\top}/\sqrt{d_{k}}\right)$, where $Q_{i}$ and $K_{j}$ are the query and key vectors of tokens $x_{i}$ and $x_{j}$, respectively, and $d_{k}$ is the key dimension. The corresponding value $V_{j}$ is then weighted by $\mathrm{Attn}_{i,j}$ to obtain the hidden representation of $x_{i}$: $z_{i}=\sum_{j}\mathrm{Attn}_{i,j}\,V_{j}$. In practice, transformers stack multiple layers and heads, producing attention weights across different levels. These attention patterns can be represented as a multi-dimensional tensor: $\mathrm{Attn}\in\mathbb{R}^{N_{l},N_{h},H,W}$, where $N_{l}$ and $N_{h}$ denote the number of layers and attention heads, respectively, and $H\times W$ represents the spatial dimensions of visual tokens, which can be reshaped into a 2D feature map. PM utilizes the $\mathrm{Attn}$ cached during forward to construct a token-level heatmap, highlighting the most relevant visual regions based on the model attention.
@@
Empirically, we find that attention maps from middle layers more accurately localize relevant objects than those from final layers. This aligns with recent findings showing that deep networks tend to aggregate and process visual information in early layers (Darcet et al. 2024), and that intermediate features can already support strong performance (Chuang et al. 2024). Hence, we construct the perception map by aggregating self-attention matrices from middle-to-late layers $l\geq\mathcal{L}$ and applying a max pooling across all heads $h$ to highlight the impact of more visually important regions. Formally, the token-level heatmap $\mathcal{H}$ can be obtained by:
@@
The basic perception map constructed above without iterative refinement could already be upsampled to guide the magnifier. As post-processing before up-sampling the token-level heatmap $\mathcal{H}$ to pixel-level perception map, we first refine the heatmap with methods similar to (Yu, Yu, and Wang 2024) to obtain a more smooth and well-distributed attention distribution for better magnification quality. Firstly, we normalize $\mathcal{H}$ by scaling it to unit interval, obtaining $\mathcal{H}^{\prime}$. Then, we amplify the variance of the normalized map by a scaling coefficient of $\alpha$, and then compress its values back into the range $[0,1]$ using a sigmoid function, resulting in $\mathcal{H}_{\text{proc}}$:
@@
Iterative Refinement
While the basic perception map already provides a simple but effective guide for magnification, it often misses some key objects’ regions when they appear multiple times. This stems from how deep vision models compress fine-grained features into a small set of tokens, commonly referred to as information registers (Darcet et al. 2024; Reddy 2024). VLMs concentrate attention to these tokens and reduce sensitivity to semantically similar but spatially distinct regions.
To improve coverage of the perception map, we introduce an iterative refinement mechanism that progressively reveals overlooked areas by suppressing highly attended tokens from prior passes. Nonetheless, directly masking key visual regions at the image level would require fine-grained localization and re-tokenization of visual tokens, making it computationally expensive. Instead, we perform attention blocking at the token level (Alg. 1, lines 4–5), which still blocks aggregated tokens, efficiently mitigates above challenges, and integrates easily into existing VLM pipelines.
At each iteration, we forward the input with an attention mask $M^{(i)}$, initialized as a vector of ones (line 1), and extract the token-level heatmap $\mathcal{H}^{(i)}$ (line 5). To identify visually dominant tokens $T^{(i)}$, we apply a simple clustering algorithm that partitions attention values into two groups (line 7), and select tokens in the one with the higher centroid as $T^{(0)}$. This leverages the observation that relevant tokens typically receive significantly higher attention than others. We then update the mask to exclude these tokens in the next iteration (line 11), encouraging the model to reallocate attention to previously under-attended regions.
The process continues until the total attention in the new heatmap drops below a threshold $\beta$ or a maximum number of iterations is reached (lines 8–10), indicating that no additional high-attention regions remain. We aggregate attention maps across all iterations and normalize them to obtain the final refined heatmap $\mathcal{H}^{*}$ (lines 13), which is then post-processed to produce the pixel-level perception map $\mathcal{P}$ (line 14). This refined map captures a broader and more accurate set of visually relevant regions, offering better guidance for magnification-enhanced decoding in downstream tasks.
Algorithm 1 Iterative Refinement on the Perception Map
1:Tokenized input $(I,Q)$, Visual Language Model (VLM), threshold $\beta$, max iterations $T_{\max}$
2:Initialize attention mask $M^{(0)}\leftarrow\mathbf{1}$
3:Initialize heatmap sum $\mathcal{H}^{*}\leftarrow\mathbf{0}$, count mask $\mathcal{C}\leftarrow\mathbf{0}$
4:for $i=0$ to $T_{\max}$ do
5:  $O^{(i)},\text{Attn}^{(i)}\leftarrow\text{VLM}(I,Q,M^{(i)})$
6:  Compute heatmap: $\mathcal{H}^{(i)}=\sum_{l\in\mathcal{L}}\max_{h}\mathrm{Attn}^{(i)}_{l,h}$
7:  Update $\mathcal{H}^{*}\leftarrow\mathcal{H}^{*}+\mathcal{H}^{(i)}$, $\mathcal{C}\leftarrow\mathcal{C}+\mathbf{1}_{M^{(i)}}$
8:  Identify $T^{(i)}\leftarrow\text{Cluster}(\mathcal{H}^{(i)})[0]$
9:  if $\sum\mathcal{H}^{(i)}<\beta$ then
10:   break
11:  end if
12:  Mask out $T^{(i)}$: $M^{(i+1)}\leftarrow M^{(i)}\cap\neg T^{(i)}$
13:end for
14:Normalize heatmap: $\mathcal{H}^{*}\leftarrow\mathcal{H}^{*}\div\mathcal{C}$
15:Post-process $\mathcal{H}^{*}$ to obtain pixel-level map $\mathcal{P}$
16:return Refined perception map $\mathcal{P}$
Attention-Based Magnification
@@
Inspired by (Zheng et al. 2019), we develop the perception magnifier as an attention-based sampling process. The perception map $\mathcal{P}$ is treated as a mass function, where regions with higher perception values are more likely to be sampled. Specifically, we first decompose $\mathcal{P}$ into marginal distributions along the horizontal and vertical dimensions and compute their cumulative distributions $\mathcal{F}_{x}(n)$ and $\mathcal{F}_{y}(n)$ as:
$\begin{split}\mathcal{F}_{x}(n):=\sum_{j=1}^{n}\max_{1\leq i\leq w}\mathcal{P}_{i,j}\quad\mathcal{F}_{y}(n):=\sum_{i=1}^{n}\max_{1\leq j\leq h}\mathcal{P}_{i,j},\end{split}$ (3)